OpenStudy (anonymous):

10. Suppose you earn $7.25 per hour working part-time for a florist. What is the minimum number of hours you must work to earn at least $125.00? Write an inequality to represent this situation and solve.

OpenStudy (anonymous):

7.25x greater than or equal to $125.00. Divide by 7.25 resulting in x greater or equal to 17.24 hours. So minimum amount of hours is 17.24 hours

OpenStudy (anonymous):

I earn $7.25 in 1 hour I earn $1 in 1/7.25 hour I earn $125 in (1/7.25)*125 = 125/7.25 hours = 17.24 hours i.e. The minimum number of hours that have to be worked is 18 hours (if u want in full hours).

OpenStudy (anonymous):

Thanks you guys